OpenStudy (anonymous):

simplify and solve for y y-2=5(x+6)

OpenStudy (solomonzelman):

\(\color{green}{ 1) }\) expand \(\color{blue}{ 5(x+6) }\) (it would be \(\color{blue}{ 5x+30 }\) \(\color{green}{ 2) }\) add \(\color{blue}{ 2 }\) to each side can you do the second step for me please?

OpenStudy (anonymous):

y - 2 = 5(x+6) y - 2 = 5x + 30 +2 +2 y=5x+32

OpenStudy (anonymous):

thank you sara019

OpenStudy (solomonzelman):

answer is always better than learning something :)